OpenStudy (anonymous):

find all x so that the distance between -2 and x is 3. A) x=5 and x=-1 B) x=-5 and x=1 C) x=5 and x=1 D) x=-5 and x=-1

OpenStudy (anonymous):

\[\left| -2-x \right|=3\]

Directrix (directrix):

A) Absolute value of (-2-x) = 3 -2-x = 3 or -2-x = -3 -x = 3+2 or -x = -3 + 2 -x = 5 or -x = -1 x = -5 or x = 1

Directrix (directrix):

B) x=-5 and x=1